# HO-Cap: A Capture System and Dataset for 3D Reconstruction and Pose Tracking of Hand-Object Interaction

### v3

@@ Abstract
Figure 1: Examples of RGB frames and the 3D shape and pose annotations of hands and objects in our dataset.
@@ 1 Introduction
An easy way to obtain pose annotations is to use motion capture (mocap) systems [4, 12]. However, mocap systems are not only expensive, but also require the use of artificial markers on hands and objects during data capture. Alternatively, several datasets, such as DexYCB [7], OakInk [51] and HOI4D [29], rely on manual labeling. Given the large number of video frames, human annotation is highly time-consuming. Recently, a few systems have been proposed to automatically or semi-automatically generate pose annotations for hands and objects [18, 24]. However, these systems are not scalable to a wide variety of objects or hand-object interactions. For example, HO-3D [18] is limited to objects with known 3D models and cannot handle unseen objects. Similarly, H2O [24] requires training an object tracker for each captured object, making it difficulty to scale to a large number of objects.
@@
Specifically, we leverage recent large pre-trained vision models, including MediaPipe [31] for hand detection and pose estimation, BundleSDF [46, 47] for 3D object reconstruction and pose estimation, FoundationPose [48] for object pose tracking, and SAM [23] and SAM2 [39] for object segmentation and tracking. To address noise and errors, we utilize multi-view consistency across eight RGB-D cameras. Additionally, we propose a Signed Distance Field (SDF)-based optimization method to refine both hand and object poses in 3D space. Consequently, our annotation pipeline can automatically process captured RGB-D videos to generate 3D shapes and poses for hands and objects. The only human annotation required is to manually select two points for each object in the first frame to generate an initial segmentation mask using SAM [23], and label the object name to register it in our database.
Using our capture system and annotation method, we created a new dataset called HO-Cap for hand-object interaction research. The dataset includes human demonstration videos of uni- and bi-manual interactions with objects, covering three interaction types: affordance-driven object use, pick-and-place, and handovers. It contains 64 videos with 656K RGB-D frames, captured from 9 subjects interacting with 64 objects. Ground-truth annotations of 3D shape and pose for hands and objects are provided for every frame. Fig. 1 shows some example images and annotations.
@@ 2 Related Work
In recent years, a number of hand-object interaction datasets have been introduced. Representative datasets are summarized in Table 1 compared to ours. In the following, we discuss in more detail the data capture and annotation process of these datasets.
Mocap vs. Natural capture. A straightforward way to obtain hand pose and object pose is to use mocap systems. By attaching reflective markers to the hands and objects, a mocap system can track these markers to obtain the hand pose and the object pose (e.g., FPHA [14], ContactPose [4], and ARCTIC [12]). However, mocap systems are costly, require calibration between mocap markers and image cameras, and introduce artifacts. In contrast, our multi-camera setup captures markerless data, eliminating the need for such calibration.
@@
Manual labeling vs. Automatic labeling. Large-scale datasets such as DexYCB [7], OakInk [51], and HOI4D [29] rely on manual labeling, which, though accurate, is labor-intensive. In contrast, some datasets use automated labeling (e.g., HO-3D [18]) or semi-automated labeling (e.g., H2O [24] and SHOWMe [43]). Fully automated methods often introduce annotation errors, while semi-automatic approaches integrate manual error correction [24] or initialize tracking processes manually [43]. Recently, HANDAL [17] introduced a semi-automatic pipeline for 6D object pose annotation, but it includes only limited dynamic human-object interactions and lacks hand annotations. In our work, we introduce a semi-automatic pipeline for annotating 3D shapes and poses of both hands and objects in videos. The only required human input is selecting two points on the object in the first frame as prompts for SAM [23] segmentation. The method then automatically annotates subsequent frames. As highlighted in Table 1, our dataset contains markerless unimanual and bimanual videos, with the capability for 3D shape reconstruction of novel objects. The most similar work to ours is H2O [24]. However, unlike H2O, our approach requires no domain-specific training for object pose trackers. Instead, we leverage pre-trained vision models with a multi-camera setup, making our annotation method more scalable across diverse hand-object interactions.
@@ 4 Annotation Method
Figure 4: Reconstructed textured meshes of the 64 objects in HO-Cap
Our annotation process begins by reconstructing 3D models of the objects. Instead of relying on pre-existing 3D meshes (e.g., YCB Object Set [5]) or 3D scanners [43], we use BundleSDF [47], a neural reconstruction method, to reconstruct a textured 3D mesh for each object, assuming that the object is rigid. BundleSDF uses a sequence of RGB-D images along with the segmentation masks of the object to precisely track its 6D pose and reconstruct a textured mesh. This enables us to use only an RGB-D camera for object reconstruction, obtaining 3D meshes for various objects, as shown in Fig. 3. To prepare the input data, we manually move and rotate an object in front of the Azure Kinect camera, ensuring exhaustive coverage of the surfaces for high-fidelity reconstruction. For segmentation, we prompt the Segmentation Anything Model (SAM) [23] with two manually selected points in the first frame to generate an accurate initial mask. This mask is then provided as input to SAM2 [39], a unified model for segmenting objects across images and videos, to track the object mask throughout the remaining frames.
Given the object masks, BundleSDF leverages feature matching based on LoFTR [42] for coarse pose initialization, followed by an online pose graph optimization to estimate the objects’ 6D poses in video frames. Simultaneously, a neural object field is trained to model object-centric geometry and appearance, refining keyframe poses to reduce tracking drift. Finally, a textured 3D mesh is extracted from the neural field using marching cubes [30] and color projection. In total, we reconstructed 64 objects in our dataset, as shown in Fig. 4. These objects are easily available for purchase online.
@@ 4.3 Hand Pose Estimation
In our initial attempts to optimize hand poses in a video sequence, we applied the same loss function used in Sec. 4.2 for object pose estimation, expecting similar results. However, practical experimentation revealed unique challenges that required a different approach for hand pose optimization. A primary issue was the difficulty in accurately segmenting the hand using the SAM and SAM2 methods, which often resulted in segmentation masks that included the forearm and other non-hand areas. This lack of precision in hand segmentation made it challenging to isolate the hand accurately for pose optimization. Additionally, we found that the higher dimensionality of the MANO hand pose parameter $\theta$, compared to the 6D pose of objects, led to overfitting when using only the SDF loss, often resulting in unrealistic hand poses.
@@ 4.4 Joint Hand-Object Pose Optimization
The point clouds of objects can be obtained using the depth images and the segmentation masks of the objects. To obtain point clouds for hands, given the optimized 3D hand keypoints, we can get the bounding box and 2D keypoints on camera images. Using these 2D keypoints along with the bounding box as input, we employ SAM [23] to generate high-quality hand masks. We further isolate hand points from the surrounding environment using a point-to-mesh distance threshold based on the initial hand pose.
@@ 4.5 HoloLens Camera Pose Estimation
4.6 Annotation Quality
We collected 64 videos featuring 9 participants performing tasks with the 64 objects in our dataset. Using our annotation method, we generated 3D shapes and poses for hands and objects. To validate the accuracy of these hand-object pose annotations, we randomly selected 800 images from 8 camera views and manually annotated visible hand joints and object keypoints. We then computed the Euclidean distance and the $R^{2}$ score [38] (coefficient of determination) between our 2D annotations and these manual annotations. The results, shown in Table 2, indicate an annotation error within 5 pixels for both hands and objects, demonstrating high precision. The $R^{2}$ Score close to 1 shows that our method captures 99.8% of the variance in the manual labels, highlighting the reliability of our annotation method. The dataset and the source code of the capture system will be released to the community. Please see the supplementary material for more details about the dataset.
Table 2: Ground-truth pose evaluation results in pixels.
Metric Object Left hand Right hand
Mean (std) $3.30(\pm 1.70)$ $4.94(\pm 3.07)$ $4.56(\pm 2.84)$
$R^{2}$ score $\uparrow$ 0.999 0.998 0.998
@@ Acknowledgement
7 MANO Shape Calibration
We calibrated the subject’s hand shape using a three-step process, assuming the left and right hands share the same shape. First, we performed initial pose estimation and 3D point collection. Starting with a rendered neutral flat MANO hand, participants are instructed to position their right hand to closely match the rendered MANO hand mesh in front of a camera. 3D points around the mesh were collected and filtered dynamically based on their distance to the mesh surface, retaining only those within a predefined threshold. The hand pose was then optimized by fitting the 3D points to the signed distance field (SDF) of the MANO model, with the shape parameter $\beta$ fixed at zero to focus solely on pose optimization. The optimized pose and filtered 3D points were saved for the next step. Second, we iteratively optimized both the hand shape $\beta$ and pose $\theta$ using the saved pose and points. This involved alternately fixing one parameter while optimizing the other to achieve better alignment between the hand mesh and the 3D point data. Finally, the optimized hand shape could be further refined by using sequences collected in the dataset. Once calibration was completed, the hand shape was fixed throughout the following optimization process for hand pose estimation.

### v4

@@ Abstract
Figure 1: Examples of RGB frames and renderings of the 3D shape and pose annotations of hands and objects in our dataset to images and in the NVIDIA Isaac Sim simulation.
@@ 1 Introduction
An easy way to obtain pose annotations is to use motion capture (mocap) systems [4, 12]. However, mocap systems are not only expensive, but also require the use of artificial markers on hands and objects during data capture. Alternatively, several datasets, such as DexYCB [7], OakInk [50] and HOI4D [29], rely on manual labeling. Given the large number of video frames, human annotation is highly time-consuming. Recently, a few systems have been proposed to automatically or semi-automatically generate pose annotations for hands and objects [18, 24]. However, these systems are not scalable to a wide variety of objects or hand-object interactions. For example, HO-3D [18] is limited to objects with known 3D models and cannot handle unseen objects. Similarly, H2O [24] requires training an object tracker for each captured object, making it difficult to scale to a large number of objects.
@@
Specifically, we leverage recent large pre-trained vision models, including MediaPipe [31] for hand detection and pose estimation, BundleSDF [45, 46] for 3D object reconstruction and pose estimation, FoundationPose [47] for object pose tracking, and SAM2 [38] for object segmentation and tracking. To address noise and errors, we utilize multi-view consistency across eight RGB-D cameras. Additionally, we propose a Signed Distance Field (SDF)-based optimization method to refine both hand and object poses in 3D space. Consequently, our annotation pipeline can automatically process captured RGB-D videos to generate 3D shapes and poses for hands and objects. The only human annotation required is to manually select two points for each object in the first frame to generate an initial segmentation mask using SAM2 [38], and label the object name to register it in our database.
Using our capture system and annotation method, we created a new dataset called HO-Cap for hand-object interaction research. The dataset includes human demonstration videos of uni- and bi-manual interactions with objects, covering three interaction types: affordance-driven object use, pick-and-place, and handovers. It contains 64 videos with 656K RGB-D frames, captured from 9 subjects interacting with 64 objects. Ground-truth annotations of 3D shape and pose for hands and objects are provided for every frame. Fig. 1 shows some examples, where we also render the annotations in the NVIDIA Isaac Sim simulator.
@@ 2 Related Work
In recent years, a number of hand-object interaction datasets have been introduced. Representative datasets are summarized in Table 1 compared to ours.
Mocap vs. Natural capture. A straightforward way to obtain hand pose and object pose is to use mocap systems. By attaching reflective markers to the hands and objects, a mocap system can track these markers to obtain the hand pose and the object pose (e.g., FPHA [14], ContactPose [4], ARCTIC [12] and OakInk2 [52]). However, mocap systems are costly, require calibration between mocap markers and image cameras, and introduce artifacts. In contrast, our multi-camera setup captures markerless data, eliminating the need for such calibration.
@@
Manual labeling vs. Automatic labeling. Large-scale datasets such as DexYCB [7], OakInk [50], and HOI4D [29] rely on manual labeling, which, though accurate, is labor-intensive. In contrast, some datasets use automated labeling (e.g., HO-3D [18]) or semi-automated labeling (e.g., H2O [24] and SHOWMe [42]). Fully automated methods often introduce annotation errors, while semi-automatic approaches integrate manual error correction [24] or initialize tracking processes manually [42]. Recently, HANDAL [17] introduced a semi-automatic pipeline for 6D object pose annotation, but it includes only limited dynamic human-object interactions and lacks hand annotations.
In our work, we introduce a semi-automatic pipeline for annotating 3D shapes and poses of both hands and objects in videos. The only required human input is selecting two points on the object in the initial frame as prompts for SAM2 [38] segmentation. The method then automatically annotates subsequent frames. As highlighted in Table 1, our dataset contains markerless unimanual and bimanual videos, with the capability for 3D shape reconstruction of novel objects. The most similar work to ours is H2O [24]. However, unlike H2O, our approach requires no domain-specific training for object pose trackers. Instead, we leverage pre-trained vision models with a multi-camera setup, making our annotation method more scalable across diverse hand-object interactions.
@@ 4 Annotation Method
Our annotation process begins by reconstructing 3D models of the objects. Instead of relying on pre-existing 3D meshes (e.g., YCB Object Set [5]) or 3D scanners [42], we use BundleSDF [46], a neural reconstruction method, to reconstruct a textured 3D mesh for each object, assuming that the object is rigid. BundleSDF uses a sequence of RGB-D images along with the segmentation masks of the object to precisely track its 6D pose and reconstruct a textured mesh. This enables us to use only an RGB-D camera for object reconstruction and obtain 3D meshes for various objects, as shown in Fig. 3.
To prepare the input data, we manually move and rotate an object in front of the Azure Kinect camera, ensuring exhaustive coverage of the surfaces for high-fidelity reconstruction. For segmentation, we prompt the SAM2 [38], a unified model for segmenting objects across images and videos, with two manually selected points in the initial frame to track the object mask throughout the remaining frames. Given the object masks, BundleSDF leverages feature matching based on LoFTR [41] for coarse pose initialization, followed by an online pose graph optimization to estimate the objects’ 6D poses in video frames. Simultaneously, a neural object field is trained to model object-centric geometry and appearance, refining keyframe poses to reduce tracking drift. Finally, a textured 3D mesh is extracted from the neural field using marching cubes [30] and color projection. In total, we reconstructed 64 objects in our dataset all of which are easily available for purchase online. For visualizations of these reconstructed objects, please refer to the supplementary material.
@@ 4.3 Hand Pose Estimation
In our initial attempts to optimize hand poses, we applied the same loss function used in Sec. 4.2 for object pose estimation, expecting similar results. However, practical experimentation revealed unique challenges that required a different approach for hand pose optimization. A primary issue was the difficulty in accurately segmenting the hand using the SAM2 method, which often resulted in segmentation masks that included the forearm and other non-hand areas. This lack of precision in hand segmentation made it challenging to isolate the hand accurately for pose optimization. Additionally, we found that the higher dimensionality of the MANO hand pose parameter $\theta$, compared to the 6D pose of objects, led to overfitting when using only the SDF loss, often resulting in unrealistic hand poses.
@@ 4.4 Joint Hand-Object Pose Optimization
The point clouds of objects can be obtained using the depth images and the segmentation masks of the objects. To obtain point clouds for hands, given the optimized 3D hand keypoints, we can get the bounding box and 2D keypoints on camera images. Using these 2D keypoints along with the bounding box as input, we employ SAM2 [38] to generate high-quality hand masks. We further isolate hand points from the surrounding environment using a point-to-mesh distance threshold based on the initial hand pose.
@@ 4.5 HoloLens Camera Pose Estimation
Table 2: Ground-truth pose evaluation results in pixel errors.
Mean (std) Object $\downarrow$ Left hand $\downarrow$ Right hand $\downarrow$
Initial Poses $3.67(\pm 3.05)$ $5.31(\pm 3.43)$ $5.16(\pm 3.73)$
SDF-Refined Poses $3.50(\pm 1.70)$ $\textbf{4.58}(\pm 2.73)$ $3.83(\pm 2.29)$
Jointly-Refined Poses $\textbf{3.42}(\pm 1.73)$ $\textbf{4.58}(\pm 2.72)$ $\textbf{3.82}(\pm 2.29)$
5 The HO-Cap Dataset
Dataset Statistics. Our HO-Cap dataset consists of 64 videos, capturing 9 participants performing 3 hand-object interaction tasks with 64 unique objects. The approximately 656K frames provide rich temporal information for studying dynamic interactions, where each frame is captured from 8 calibrated RealSense cameras plus a first-person-view camera from the HoloLens, facilitating 3D reconstruction and egocentric understanding. The 64 different objects, each with a textured 3D mesh model, enabling fine-grained 6D pose annotations, and a diverse set of 9 subjects ensures variability in hand shapes, grasping styles, and interaction patterns. The both single-hand and bimanual interactions, supporting tasks that require complex hand coordination. Table 3 provides handedness statistics for left/right hands and bi-manual interactions, and a breakdown of task categories. More dtails about the object shape and hand-object pose diversity are provided in supplementary material.
Annotation Quality. Using our annotation method, we generated 3D shapes and world space poses for hands and objects. To validate the annotation accuracy, we randomly selected 800 images across 8 RealSense camera views and manually annotated visible hand joints and object keypoints. The object keypoints were chosen from predefined mesh vertices that are easily identifiable. We then computed the Euclidean distance between the 2D projections of our 3D annotations and these human labeled 2D points. Table 2 presents the evaluation results. 1) The final annotation error is within 5 pixels for both hands and objects, demonstrating the reliability of our method, and (2) The error consistently decreases across annotation refinement steps, validating the effectiveness of our optimization strategy.
Table 3: Detailed Dataset Statistics. The statistics are grouped by handness (Right, Left, Both) and tasks (T1: pick-and-place, T2: handover between hands, T3: affordance usage).
Statistics Handness Tasks
R L B T1 T2 T3
#sequence 35 8 21 28 21 15
#frames 41,327 8,373 23,244 29,706 23,244 19,994
@@ Acknowledgement
8 Details on Data Collection
8.1 MANO Shape Calibration
We calibrated each subject’s hand shape using a three-step process, assuming both hands share the same shape. First, we performed initial pose estimation and 3D point collection. Participants were instructed to align their right hand with a rendered neutral MANO hand mesh displayed in front of a camera (Fig. 6). During this process, 3D points around the mesh were dynamically collected and filtered based on their distance to the mesh surface, retaining only those within a predefined threshold. The hand pose was then optimized by fitting the 3D points to the signed distance field (SDF) of the MANO model, keeping the shape parameter $\beta$ fixed at zero to focus solely on pose refinement. The optimized pose and filtered 3D points were saved for the next step. Second, we iteratively optimized both the hand shape $\beta$ and pose $\theta$ using the saved pose and points. This involved alternately fixing one parameter while optimizing the other, gradually improving alignment between the hand mesh and the collected 3D point data. Finally, the optimized hand shape could be further refined by using sequences collected in the dataset. Once finalized, the calibrated hand shape was fixed throughout the following optimization process for hand pose estimation.
Figure 6: Visualization of the rendered MANO hand mesh used as a reference for participants during calibration.
8.2 Synchronization of the Camera Streams
To facilitate efficient control and data collection, all cameras are integrated into a ROS server for seamless acquisition. The 8 static RealSense cameras are wired to the server, where they publish synchronized RGB and depth images. The HoloLens, connected via TCP/IP, operates in Research Mode to access and publish RGB camera streams and head tracking data. All image frames and pose data are recorded into a single ROS bag file, with synchronization performed based on timestamps to ensure temporal alignment.
9 Properties of HO-Cap
9.1 Annotation Details
The HO-Cap dataset provides MANO-based 3D hand pose and 6D object pose annotations, optimized within a global world frame using 8 RealSense cameras. Camera intrinsics and extrinsics for all views are included, enabling the transformation of world-frame annotations into camera-frame 3D poses. The 2D hand joint keypoints are obtained by projecting the 3D hand joints onto the image plane. Additionally, segmentation masks offer pixel-wise annotations for objects and hands, facilitating precise scene understanding. The First-Person View (FPV) data from the HoloLens headset provides egocentric perspectives crucial for human perception research and assistive AR applications.
9.2 Visualization and Scene Simulation
The HO-Cap dataset provides textured 3D meshes for 64 objects (Fig. 8), compatible with physics simulators, enabling realistic scene reconstruction and interaction modeling. For sequence data visualization and scene replay in Isaac Sim, please refer to the supplementary video.
9.3 Shape and Pose Diversity
As illustrated in Fig. 7, our dataset (1) encompasses a diverse range of hand-held object shapes and sizes. and (2) offers a broader variety of hand poses compared to HO-3D [18], capturing more natural and dynamic interactions.
Figure 7: Distribution of (a) object volume ($mm^{3}$), (b) hand pose with comparison to HO-3D (first two MANO PCA coefficients).
10 Comparison to Similar Datasets
The most related datasets to HO-Cap are DexYCB[7], HO-3D[18], and H2O [24]. DexYCB focuses on 20 objects from the YCB benchmark, primarily capturing single-hand grasping and handover interactions using 8 camera views. In contrast, HO-Cap provides 64 unique objects with textured 3D meshes, offering greater object diversity. Additionally, HO-Cap includes both single-hand and bimanual interactions and incorporates egocentric (FPV) data, enhancing its applicability for AR/VR and human-centered AI research. HO-3D features grasping-centric hand poses with 10 objects, captured from 1–5 views per sequence, and relies on marker-based motion capture for annotations. In contrast, HO-Cap supports a significantly wider range of hand poses and grasps across 64 objects and employs a semi-automatic optimization-based annotation pipeline that does not require markers. H2O is one of the few datasets that support bimanual hand-object interactions while providing 6D object poses and 3D hand poses. However, it is limited to 8 objects. In comparison, HO-Cap extends this to 64 objects, enabling a wider range of bimanual grasping scenarios and more diverse hand-object interactions.